Assume the length X, in minutes, of a particular type of telephone conversation is a random variable with probability density function

gPNq7G+KgStuwAAAAASUVORK5CYII=

(a) Determine the mean length E(X) of this type of telephone conversation.

(b) Find the variance and standard deviation of X.

(c) Find E[(X+4)2]i2HyUzHDZnwAAAABJRU5ErkJggg== .
Answer #1

Sol Given that Assume the length x, in minutes of a particular type of telephone conversation is a gandom Vasiable with probaWe know that • r(n = rent dx ,nxo (n+1) = no if nis integer From equation o Elx) = 54(2) E(x) = 4X!! F(x) = 4x1 JE(x) = 4 b)T08 = Get 8h = +x8 +93 +6€ = (*)28 +91+(x)2 = [x8+21x] = = E[(x+4)%] = [ *4+ 2x4 xx] () To find E [(x+4)*] :-